HELP! I'm an old dog learning new stuff. I'm having trouble getting this to click in my old brain. I'm supposed to figure this out by using dimensional analysis.... Thanks so very much for your help!
If a block of ice measures 24 in. x 36 in. x 18 in.
(a) What is the volume of this block in cubic meters?
(b) What is the total surface area of the block in square centimeters?

Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

If a block of ice measures 24 in. x 36 in. x 18 in.
----------------
1 inch is 2.54 cetimeters
1 cm = 0.01 meter
------------------------------------
24 in = 24*2.54 cm = 60.96 cm = 0.6096 meters
36 in = 36*2.54 cm = 91.44 cm = 0.9144 meters
18 in = 18*2.54 cm = 45.72 cm = 0.4572 meters
=---------------------------------------
(a) What is the volume of this block in cubic meters?
Volume = the product of the meter measures = 0.25485 cu meters
-----------------------------------------------------------------------
(b) What is the total surface area of the block in square centimeters?
surface area = 2(60.96*91.44) + 2(60.96*45.72) + 2(91.44*45.72)
= 11148.3648 + 5574.1924 + 8361.2736
= 25083.8308 sq. cm.
=============================
Cheers,
Stan H.

Answer by richwmiller(17219)

You need your conversion factor to convert from cubic inches to cubic meters and square inches to square centimeters
1 inches = 0.0254 meters
http://vulcan.wr.usgs.gov/Miscellaneous/ConversionTables/conversion_table.html
you can convert before multiplying or after